A car, starting from rest at position X, accelerates up a hill. The car reaches a speed of 10m / s at position Y. The kinetic energy of the car at position Y is equal to its gain in gravitational potential energy from X to Y. X Y gain in height of car Take the gravitational field strength g to be 10N/ kg. What is the gain in height of the car between X and Y?


1
Expert's answer
2021-02-22T10:23:23-0500

The kinetic energy of the car at position Y is the following:


"K = \dfrac{mv^2}{2}"

where "m" is the mass of the car, and "v = 10 m\/s" is its speed. The gravitational potential energy from X to Y, which is given as follows:


"U = mgh"

where "g = 10 N\/kg" is the gravitational accleration, and "h" is the gain in height of the car between X and Y.

Since these two quantities are equal, obtain:


"\dfrac{mv^2}{2} = mgh\\
h = \dfrac{v^2}{2g}\\
h = \dfrac{10^2}{2\cdot 10} = 5 m"

Answer. 5 m.